OpenStudy (anonymous):

1. Let [x] denote the integer part of x. ( eg. [2.56788] = 2 ) Prove that [(2 + √3)^n] is an odd number, where n is a non-negative integer. 2. Let {x} denote the decimal part of x. ( eg. {2.44598} = 0.44598 ) If for natural number n , R = (2 + √3)^n and f = {R} , Prove that R(1 - f) = 1

OpenStudy (anonymous):

Do you need a hit?

OpenStudy (anonymous):

not yet

OpenStudy (experimentx):

lol ... we don't need 'hit'

OpenStudy (anonymous):

haha hint

OpenStudy (experimentx):

where do you get your problems ... most of them are insanely difficult

OpenStudy (anonymous):

internet

OpenStudy (anonymous):

and Hong Kong Advanced Level Examination (Pure Math)

OpenStudy (anonymous):

These two questions are not :)

OpenStudy (anonymous):

Okay. I got it. Should I post my solution or do guys want to try it too?

OpenStudy (anonymous):

can't wait

OpenStudy (anonymous):

Haha. Okay. Brb.

OpenStudy (anonymous):

so sid bring ur answer

OpenStudy (anonymous):

For convinience I type 3^1/2 as k. So, Let (2 + k) ^ n be some I + f1 ---- Where I is the integral part and f the fractional.  And Let ( 2 - k)^n be some f2 ( It has to be fractional) Adding the two we get, I1 + + f1 + f2 = ( Something which is integral, As on adding the irrational parts cancel out to give only an integer.) Therefore, f1 + f2 + I1is an integer. => f1 + f2 is an integer. Now 0<f1<1 and f2 also has the same range. Therefore 0<f1+f2<2. For it to be an integer f1 + f2 = 1. Therefor I1 + 1 = even. ( Its Twice of something) Therefore I1 = Greatest integer of given expression is odd.